# Quantum DeepONet: Neural operators accelerated by quantum computing

### v1

@@ Abstract
In the realm of computational science and engineering, constructing models that reflect real-world phenomena requires solving partial differential equations (PDEs) with different conditions. Recent advancements in neural operators, such as deep operator network (DeepONet), which learn mappings between infinite-dimensional function spaces, promise efficient computation of PDE solutions for a new condition in a single forward pass. However, classical DeepONet entails quadratic complexity concerning input dimensions during evaluation. Given the progress in quantum algorithms and hardware, here we propose to utilize quantum computing to accelerate DeepONet evaluations, yielding complexity that is linear in input dimensions. Our proposed quantum DeepONet integrates unary encoding and orthogonal quantum layers. We benchmark our quantum DeepONet using a variety of PDEs, including the antiderivative operator, advection equation, and Burgers’ equation. We demonstrate the method’s efficacy in both ideal and noisy conditions. Furthermore, we show that our quantum DeepONet can also be informed by physics, minimizing its reliance on extensive data collection. Quantum DeepONet will be particularly advantageous in applications in outer loop problems which require to explore parameter space and solving the corresponding PDEs, such as uncertainty quantification and optimal experimental design.
@@ 1 Introduction
To address these limitations, deep neural operators have gained popularity for learning the mapping between infinite-dimensional spaces of functions through data [13, 14, 15, 16, 17, 18]. Once trained, neural operators are able to efficiently evaluate the PDE solutions for a new PDE instance in a single forward pass. Additionally, the output of neural operators can be discretized at different levels of resolutions or evaluated at any points. The training of neural operators can also incorporate physics priors [19, 20], aligning the concept of PINNs, which has been shown to enhance accuracy significantly. The main categories of neural operators include integral kernel operators [15, 14, 21], transformer-based neural operators [17, 16], and DeepONet [13]. Integral kernel operators, such as Fourier neural operator (FNO) [14], leverage iterative learnable kernel integration, but are usually restricted to grids. Transformer-based neural operator has larger model capacity, but relies on sufficient data to achieve optimal performance. DeepONet, grounded in universal approximation theorem [22], on the other hand, can evaluate the solution of PDEs at any points in a mesh-free manner. There have been a wide range of developments of DeepONet [23, 24, 25, 26, 27], highlighting its adaptability in various complex systems.
@@
Neural operators present an ideal application scenario for quantum neural networks designed for accelerating the evaluation process, especially in situations where they are evaluated repeatedly in “outer-loop problems”, such as forward uncertainty propagation and optimal experimental design. There is a recent development of quantum Fourier neural operator (QFNO) [40]. Utilizing a new form of the quantum Fourier transform, QFNO is expected to be substantially faster than classical FNO in evaluation. Instead of the linear number of evaluations required by classical FNO, QFNO only needs a logarithmic number of evaluations of the initial condition function, offering a significant improvement in efficiency. The success of QFNO motivates us to explore the possibility of accelerating other neural operators, such as DeepONet.
@@
The paper is organized as follows. We first present the algorithm and architecture of quantum DeepONet in Section 2. In Section 3, we illustrate the ideal quantum simulation results of different applications of our quantum DeepONet. Then we investigate quantum noise and show the performance of the quantum DeepONet under two different noise models in Section 4. Finally, we conclude our work and discuss the limitations in Section 5. The background concepts related to quantum computing is provided in Appendices A and B.
@@ 2 Methods
In this section, we first introduce a specific quantum circuit for network layers in Section 2.1, referred to as “quantum layers,” which are designed for constructing quantum orthogonal neural networks (Section 2.2). Building on these foundations, we propose a novel quantum DeepONet structure by synthesizing multiple quantum layers in Section 2.3. The training method and loss function are detailed in Section 2.4. Furthermore, in addition to data driven training, we also propose to use physics-informed loss function, developing quantum physics-informed DeepONet (QPI-DeepONet) in Section 2.5.
@@ 2.1 Quantum methods for network layers
A classical neural network layer, with the input $\mathbf{x}\in\mathbb{R}^{n}$ and output $\mathbf{x^{\prime}}\in\mathbb{R}^{m}$, takes the form $\mathbf{x^{\prime}}=\sigma(\mathbf{W}\mathbf{x}+\mathbf{b})$. Here, $\mathbf{W}\in\mathbb{R}^{m\times n}$ represents the weight matrix, $\mathbf{b}\in\mathbb{R}^{m}$ is the bias, and $\sigma$ is the activation function. As demonstrated by Ref. [46], the matrix multiplication $\mathbf{W}\mathbf{x}$ can be accelerated by substituting the classical matrix multiplication with quantum matrix multiplication. The neural network layer accelerated by this quantum algorithm is referred as a quantum layer. We provide a detailed explanation of each step of a “quantum layer”, beginning with an introduction to the basic gate, the reconfigurable beam splitter (RBS) gate, used in our method (Section 2.1.1). The whole process involves three key steps to handle classical data on a quantum computer: (1) loading the classical data onto the quantum circuit (Section 2.1.2), (2) performing matrix multiplication on quantum computer (Section 2.1.3), and (3) converting the resulting quantum data back into classical data (Section 2.1.4). We summarize and provide the complexity of each step in Section 2.1.5.
Figure 1: The circuits of a quantum layer. A quantum layer is composed of data loading, pyramidal circuit, and tomography. An ancillary qubit is included for the purpose of tomography. The vertical lines represent the two-qubit RBS gates, while the $\theta_{1}$ and $\alpha_{i}$ correspond to the parameter of the gate. We provide the example of data loader for loading the classical vector $\mathbf{x}\in\mathbb{R}^{4}$ with $\|\mathbf{x}\|_{2}=1$. We demonstrate quantum pyramidal circuit using all of the seven examples. $\mathbf{W}\in\mathbb{R}^{4\times 4}$, $\mathbf{W}\in\mathbb{R}^{4\times 3}$, and $\mathbf{W}\in\mathbb{R}^{3\times 4}$ share the same pyramidal circuit. The following circuit are other examples of $m\neq n$ cases: $\mathbf{W}\in\mathbb{R}^{4\times 2}$, $\mathbf{W}\in\mathbb{R}^{4\times 1}$, $\mathbf{W}\in\mathbb{R}^{2\times 4}$ and $\mathbf{W}\in\mathbb{R}^{1\times 4}$.
@@ 2.1.2 Loading classical data input
To load vector $\mathbf{x}$ on to a quantum state, it is crucial to ensure that the norm $\|\mathbf{x}\|_{2}=1$, as required by the probabilistic nature of quantum mechanics. If $\|\mathbf{x}\|_{2}\neq 1$, normalization is required. We append an additional dimension to $\mathbf{x}$ at the first quantum layer in the neural network, which keeps the norm of $\mathbf{x}$ at 1 and in the meantime store the information of the original norm of $\mathbf{x}$. Specifically, each element of $\mathbf{x}$ is first rescaled to the range $[-1,1]$. Then the value $\sqrt{1-\sum_{i}x_{i}^{2}/d}$ is assigned to the new dimension, where $d$ represent the original dimensionality of $\mathbf{x}$. This procedure can be viewed as data preprocessing before training, transforming the original $\mathbf{x}$ into
$\begin{pmatrix}x_{1}\\ x_{2}\\ \ldots\\ x_{m}\\ \sqrt{1-\sum_{i}x_{i}^{2}/d}\end{pmatrix},$
where $x_{i}$ is the $i$th element of $\mathbf{x}$. For subsequent quantum layers in the neural network, we simply dividing $\mathbf{x}$ by $\|\mathbf{x}\|_{2}$ before loading the data.
The circuit for loading data is shown in Fig. 1 bottom. If $m=n$, the quantum circuit we adopt will have $n$ qubits, initialized such that the first qubit is at state |$\ket{1}$, while remaining qubits are |$\ket{0}$. Then we apply a series of RBS gates parameterized by $(\alpha_{1},\alpha_{2},\ldots,\alpha_{n-1})$, where
@@
The configuration, where one qubit is in state |$\ket{1}$, while all others are in state |$\ket{0}$, is referred to as a “unary state.” For simplicity, the $j$th unary state is denoted as |$\ket{\mathbf{e}_{j}}$. Therefore, the information of $\mathbf{x}$ is encapsulated in |$\ket{x}$ represented as the superposition of these unary states. Once data is loaded into the superposition of unary states, all of our subsequent operations, which utilize the RBS gate and only include transformations between |$\ket{\mathbf{e}_{j}}$ states, are effectively confined to these unary states. This implies that the unary subspace throughout entire process, allowing us to employ the tomography, which will be further explained in Section. 2.1.4.
On the other hand, when output and input dimensions are different $(m\neq n)$, the number of qubits required in the circuit will be $\max(m,n)$. If output dimension is smaller than input dimension, i.e., $m<n$, the data is loaded onto all of the $n$ qubits. Conversely, if $m>n$, the classical vector $\mathbf{x}$ is loaded on to the bottom $n$ qubits in the circuit, leaving upper $m-n$ qubits at |$\ket{0}$.
@@ 2.1.3 Quantum pyramidal circuit
We first introduce the quantum pyramidal circuit for $m=n$ cases. The basic idea of this method is to decomposes the orthogonal matrix $\mathbf{W}$ into a series of rotation matrices, which can be represented by RBS gates. These decomposed rotation matrices can be parameterized with angles ${\theta_{1},\theta_{2}\ldots,\theta_{d}}$, where $d=n(n-1)/2$. All of the parameterized RBS gates are arranged in a pyramid configuration. We take the $\mathbf{W}\in\mathbb{R}^{4\times 4}$ matrix in Fig. 1 as an example. This circuit conducts the following operation on the loaded vector $x$:
$\mathbf{y}=\underbrace{\begin{pmatrix}C_{\theta_{1}}&S_{\theta_{1}}&&\\ -S_{\theta_{1}}&C_{\theta_{1}}&&\\ &&1&\\ &&&1\end{pmatrix}\begin{pmatrix}1&&&\\ &C_{\theta_{2}}&S_{\theta_{2}}&\\ &-S_{\theta_{2}}&C_{\theta_{2}}&\\ &&&1\end{pmatrix}\begin{pmatrix}C_{\theta_{3}}&S_{\theta_{3}}&&\\ -S_{\theta_{3}}&C_{\theta_{3}}&&\\ &&C_{\theta_{4}}&S_{\theta_{4}}\\ &&-S_{\theta_{4}}&C_{\theta_{4}}\end{pmatrix}\begin{pmatrix}1&&&\\ &C_{\theta_{5}}&S_{\theta_{5}}&\\ &-S_{\theta_{5}}&C_{\theta_{5}}&\\ &&&1\end{pmatrix}\begin{pmatrix}C_{\theta_{6}}&S_{\theta_{6}}&&\\ -S_{\theta_{6}}&C_{\theta_{6}}&&\\ &&1&\\ &&&1\end{pmatrix}}_{\mathbf{W}}\mathbf{x},$
where $C_{\theta_{j}}$ and $S_{\theta_{j}}$ are $\cos\theta_{j}$ and $\sin\theta_{j}$ for any $j$, respectively. Therefore, the resulting quantum state is
|$\ket{y}=\ket{Wx}=\sum_{ij}W_{ij}x_{i}\ket{\mathbf{e}_{j}}.$
If $m\neq n$, the construction of pyramidal circuit is the same as Ref. [46]. Examples of this include $\mathbf{W}\in\mathbb{R}^{4\times 1}$, $\mathbf{W}\in\mathbb{R}^{1\times 4}$ and so on, as shown in Fig. 1. Note that for $|m-n|=1$ cases, the pyramidal circuit is the same as $m=n$ cases. However, due to the difference in data loading and tomography process, the quantum layer is actually distinct.
@@ 2.1.4 Tomography for extracting classical output
After performing matrix multiplication in quantum space, it is necessary to convert the quantum information back to classical form for further processing, such as adding bias and applying non-linear transformation. This process is known as tomography. Tomography could commonly be expensive in terms of quantum resources when extract the complete information from quantum states [47, 48, 49]. However, in our method, the usage of unary state sparsely encodes information in Hilbert space and provides a feasible and efficient tomography method. This tomography method, proposed by Ref. [46], is illustrated in Fig. 1 middle.
@@
We introduce an ancillary qubit and implement a Hadamard ($H$) and a CNOT gate between the ancillary qubit and the first data loader qubit before loading data (see Appendix A for the definition of gates). After the pyramid gate, the circuit performs an adjoint operation of the data loader of a uniform norm-1 vector $\left(\frac{1}{\sqrt{r}},\frac{1}{\sqrt{r}},\ldots,\frac{1}{\sqrt{r}}\right)$, where $r=\max(m,n)$ represents the number of qubits excluding the ancillary qubit. This is followed by an X gate and CNOT gate.
@@
Finally, we load $\left(\frac{1}{\sqrt{r}},\frac{1}{\sqrt{r}},\ldots,\frac{1}{\sqrt{r}}\right)$ and a Hadamard gate. In this way, the output can be represented by
$y_{j}=\sum_{i}W_{ij}x_{i}=\sqrt{r}({\rm Pr}[0,\mathbf{e}_{j}]-{\rm Pr}[1,\mathbf{e}_{j}]),$ (1)
where ${\rm Pr}[\xi,\mathbf{e}_{j}]$ means the the ancillary qubit is measured as a classical bit $\xi$, for $\xi\in\{0,1\}$, and the rest qubits are measured as $\mathbf{e}_{j}$. As a result, the value of $\sum_{i}W_{ij}x_{i}$ can be simply computed from the probabilities of |$\ket{0,\mathbf{e}_{j}}$ and |$\ket{1,\mathbf{e}_{j}}$ for all needed $j$.
@@
When the output dimension is smaller than the input dimension ($m<n$), the tomography circuit is still the same, but only the information of bottom $m$ qubits are finally considered. In other words, the $\mathbf{e}_{j}$ in Eq. (1) refers to $j$th unary state for the bottom $m$ qubits. Consequently, the output $\sum_{i}W_{ij}x_{i}$ is restricted to size $m$.

### v2

@@ Abstract
In the realm of computational science and engineering, constructing models that reflect real-world phenomena requires solving partial differential equations (PDEs) with different conditions. Recent advancements in neural operators, such as deep operator network (DeepONet), which learn mappings between infinite-dimensional function spaces, promise efficient computation of PDE solutions for a new condition in a single forward pass. However, classical DeepONet entails quadratic complexity concerning input dimensions during evaluation. Given the progress in quantum algorithms and hardware, here we propose to utilize quantum computing to accelerate DeepONet evaluations, yielding complexity that is linear in input dimensions. Our proposed quantum DeepONet integrates unary encoding and orthogonal quantum layers. We benchmark our quantum DeepONet using a variety of PDEs, including the antiderivative operator, advection equation, and Burgers’ equation. We demonstrate the method’s efficacy in both ideal and noisy conditions. Furthermore, we show that our quantum DeepONet can also be informed by physics, minimizing its reliance on extensive data collection. Quantum DeepONet will be particularly advantageous in applications in outer loop problems which require exploring parameter space and solving the corresponding PDEs, such as uncertainty quantification and optimal experimental design.
@@ 1 Introduction
To address these limitations, deep neural operators have gained popularity for learning the mapping between infinite-dimensional spaces of functions through data [13, 14, 15, 16, 17, 18]. Once trained, neural operators are able to efficiently evaluate the PDE solutions for a new PDE instance in a single forward pass. Additionally, the output of neural operators can be discretized at different levels of resolutions or evaluated at any point. The training of neural operators can also incorporate physics priors [19, 20], aligning the concept of PINNs, which has been shown to enhance accuracy significantly. The main categories of neural operators include integral kernel operators [15, 14, 21], transformer-based neural operators [17, 16], and DeepONet [13]. Integral kernel operators, such as Fourier neural operator (FNO) [14], leverage iterative learnable kernel integration, but are usually restricted to grids. Transformer-based neural operator has larger model capacity but relies on sufficient data to achieve optimal performance. DeepONet, grounded in universal approximation theorem [22], on the other hand, can evaluate the solution of PDEs at any point in a mesh-free manner. There has been a wide range of developments of DeepONet [23, 24, 25, 26, 27], highlighting its adaptability in various complex systems.
@@
Neural operators present an ideal application scenario for quantum neural networks designed to accelerate the evaluation process, especially in situations where they are evaluated repeatedly in “outer-loop problems”, such as forward uncertainty propagation and optimal experimental design. There is a recent development of quantum Fourier neural operator (QFNO) [40]. Utilizing a new form of the quantum Fourier transform, QFNO is expected to be substantially faster than classical FNO in evaluation. Instead of the linear number of evaluations required by classical FNO, QFNO only needs a logarithmic number of evaluations of the initial condition function, offering a significant improvement in efficiency. The success of QFNO motivates us to explore the possibility of accelerating other neural operators, such as DeepONet.
@@
The paper is organized as follows. We first present the algorithm and architecture of quantum DeepONet in Section 2. In Section 3, we illustrate the ideal quantum simulation results of different applications of our quantum DeepONet. Then we investigate quantum noise and show the performance of the quantum DeepONet under two different noise models in Section 4. Finally, we conclude our work and discuss the limitations in Section 5. The background concepts related to quantum computing are provided in Appendices A and B.
@@ 2 Methods
In this section, we first introduce a specific quantum circuit for network layers in Section 2.1, referred to as “quantum layers”, which are designed for constructing quantum orthogonal neural networks (Section 2.2). Building on these foundations, we propose a novel quantum DeepONet structure by synthesizing multiple quantum layers in Section 2.3. The training method and loss function are detailed in Section 2.4. Furthermore, in addition to data-driven training, we also propose to use a physics-informed loss function, developing quantum physics-informed DeepONet (QPI-DeepONet) in Section 2.5.
@@ 2.1 Quantum methods for network layers
A classical neural network layer, with the input $\mathbf{x}\in\mathbb{R}^{n}$ and output $\mathbf{x^{\prime}}\in\mathbb{R}^{m}$, takes the form $\mathbf{x^{\prime}}=\sigma(\mathbf{W}\mathbf{x}+\mathbf{b})$. Here, $\mathbf{W}\in\mathbb{R}^{m\times n}$ represents the weight matrix, $\mathbf{b}\in\mathbb{R}^{m}$ is the bias, and $\sigma$ is the activation function. As demonstrated by Ref. [46], the matrix multiplication $\mathbf{W}\mathbf{x}$ can be accelerated by substituting the classical matrix multiplication with quantum matrix multiplication. The neural network layer accelerated by this quantum algorithm is referred as a quantum layer. We provide a detailed explanation of each step of a “quantum layer”, beginning with an introduction to the basic gate, the reconfigurable beam splitter (RBS) gate, used in our method (Section 2.1.1). Because adding bias and applying non-linear transformation still require classical computation, the whole quantum layer, as depicted in Fig. 1, involves three key steps to handle classical data on a quantum computer: (1) loading the classical data onto the quantum circuit (Section 2.1.2), (2) performing matrix multiplication on quantum computer (Section 2.1.3), and (3) converting the resulting quantum data back into classical data (Section 2.1.4). In Section 2.1.5, we illustrate a specific error mitigation method resulting from the unary encoding. Finally, we summarize and provide the complexity of each step in Section 2.1.6.
Figure 1: The circuits of a quantum layer. A quantum layer is composed of data loading, a pyramidal circuit, and tomography. An ancillary qubit is included for tomography. The vertical lines represent the two-qubit RBS gates, while the $\theta_{i}$ and $\alpha_{i}$ correspond to the parameter of the gate. We provide the example of a data loader for loading the classical vector $\mathbf{x}\in\mathbb{R}^{4}$ with $\|\mathbf{x}\|_{2}=1$. We demonstrate a quantum pyramidal circuit using all of the seven examples. $\mathbf{W}\in\mathbb{R}^{4\times 4}$, $\mathbf{W}\in\mathbb{R}^{4\times 3}$, and $\mathbf{W}\in\mathbb{R}^{3\times 4}$ share the same pyramidal circuit. The following circuit are other examples of $m\neq n$ cases: $\mathbf{W}\in\mathbb{R}^{4\times 2}$, $\mathbf{W}\in\mathbb{R}^{4\times 1}$, $\mathbf{W}\in\mathbb{R}^{2\times 4}$ and $\mathbf{W}\in\mathbb{R}^{1\times 4}$.
@@ 2.1.2 Loading classical data input
To load vector $\mathbf{x}$ onto a quantum state, it is crucial to ensure that the norm $\|\mathbf{x}\|_{2}=1$, as required by the probabilistic nature of quantum mechanics. If $\|\mathbf{x}\|_{2}\neq 1$, normalization is required. We append an additional dimension to $\mathbf{x}$ at the first quantum layer in the neural network, which keeps the norm of $\mathbf{x}$ at 1 and in the meantime stores the information of the original norm of $\mathbf{x}$. Specifically, each element of $\mathbf{x}$ is first rescaled to the range $[-1,1]$. Then the value $\sqrt{1-\sum_{i}x_{i}^{2}/n}$ is assigned to the new dimension. This procedure can be viewed as data preprocessing before training, transforming the original $\mathbf{x}$ into
$\begin{pmatrix}x_{1}\\ x_{2}\\ \ldots\\ x_{n}\\ \sqrt{1-\sum_{i}x_{i}^{2}/n}\end{pmatrix},$
where $x_{i}$ is the $i$th element of $\mathbf{x}$. For subsequent quantum layers in the neural network, we simply divide $\mathbf{x}$ by $\|\mathbf{x}\|_{2}$ before loading the data.
The circuit for loading data is shown in Fig. 1 bottom. If $m=n$, the quantum circuit we adopt will have $n$ qubits, initialized such that the first qubit is at state |$\ket{1}$ while the remaining qubits are |$\ket{0}$. Then we apply a series of RBS gates parameterized by $(\alpha_{1},\alpha_{2},\ldots,\alpha_{n-1})$, where
@@
The configuration, where one qubit is in state |$\ket{1}$, while all others are in state |$\ket{0}$, is referred to as a “unary state”. For simplicity, the $j$th unary state is denoted as |$\ket{\mathbf{e}_{j}}$. Therefore, the information of $\mathbf{x}$ is encapsulated in |$\ket{x}$ represented as the superposition of these unary states. Once data is loaded into the superposition of unary states, all of our subsequent operations, which utilize the RBS gate and only include transformations between |$\ket{\mathbf{e}_{j}}$ states, are effectively confined to these unary states. This implies that the unary subspace throughout the entire process, allows us to employ the tomography, which will be further explained in Section 2.1.4.
On the other hand, when output and input dimensions are different $(m\neq n)$, the number of qubits required in the circuit will be $\max(m,n)$. If the output dimension is smaller than the input dimension, i.e., $m<n$, the data is loaded onto all $n$ qubits. Conversely, if $m>n$, the classical vector $\mathbf{x}$ is loaded onto the bottom $n$ qubits in the circuit, leaving upper $m-n$ qubits at |$\ket{0}$.
@@ 2.1.3 Quantum pyramidal circuit
We first introduce the quantum pyramidal circuit for $m=n$ cases. The basic idea of this method is to decompose the orthogonal matrix $\mathbf{W}$ into a series of rotation matrices, which can be represented by RBS gates. These decomposed rotation matrices can be parameterized with angles ${\theta_{1},\theta_{2}\ldots,\theta_{d}}$, where $d=n(n-1)/2$. All of the parameterized RBS gates are arranged in a pyramid configuration. We take the $\mathbf{W}\in\mathbb{R}^{4\times 4}$ matrix in Fig. 1 as an example. On the loaded vector $\mathbf{x}$, the pyramidal portion of the circuit conducts $\mathbf{y}=\mathbf{W}\mathbf{x}$, where $\mathbf{W}$ is
$\begin{pmatrix}C_{\theta_{1}}&S_{\theta_{1}}&&\\ -S_{\theta_{1}}&C_{\theta_{1}}&&\\ &&1&\\ &&&1\end{pmatrix}\begin{pmatrix}1&&&\\ &C_{\theta_{2}}&S_{\theta_{2}}&\\ &-S_{\theta_{2}}&C_{\theta_{2}}&\\ &&&1\end{pmatrix}\begin{pmatrix}C_{\theta_{3}}&S_{\theta_{3}}&&\\ -S_{\theta_{3}}&C_{\theta_{3}}&&\\ &&C_{\theta_{4}}&S_{\theta_{4}}\\ &&-S_{\theta_{4}}&C_{\theta_{4}}\end{pmatrix}\begin{pmatrix}1&&&\\ &C_{\theta_{5}}&S_{\theta_{5}}&\\ &-S_{\theta_{5}}&C_{\theta_{5}}&\\ &&&1\end{pmatrix}\begin{pmatrix}C_{\theta_{6}}&S_{\theta_{6}}&&\\ -S_{\theta_{6}}&C_{\theta_{6}}&&\\ &&1&\\ &&&1\end{pmatrix}.$
$C_{\theta_{j}}$ and $S_{\theta_{j}}$ are $\cos\theta_{j}$ and $\sin\theta_{j}$ for any $j$, respectively. Therefore, the resulting quantum state is
|$\ket{y}=\ket{Wx}=\sum_{i,j}W_{ji}x_{i}\ket{\mathbf{e}_{j}}.$
If $m\neq n$, the construction of the pyramidal circuit is the same as Ref. [46]. Examples of this include $\mathbf{W}\in\mathbb{R}^{4\times 1}$, $\mathbf{W}\in\mathbb{R}^{1\times 4}$ and so on, as shown in Fig. 1. Note that for $|m-n|=1$ cases, the pyramidal circuit is the same as $m=n$ cases. However, due to the difference in data loading and tomography process, the quantum layer is distinct.
@@ 2.1.4 Tomography for extracting classical output
After performing matrix multiplication in quantum space, it is necessary to convert the quantum information back to classical form for further processing, such as adding bias and applying non-linear transformation. This process is known as tomography. Tomography could commonly be expensive in terms of quantum resources when extracting the complete information from quantum states [47, 48, 49]. However, in our method, the usage of unary state sparsely encodes information in Hilbert space and provides a feasible and efficient tomography method. This tomography method, proposed by Ref. [46], is illustrated in Fig. 1 middle.
@@
We introduce an ancillary qubit and implement a Hadamard ($H$) and a CNOT gate between the ancillary qubit and the first data loader qubit before loading data (see Appendix A for the definition of gates). After the pyramid gate, the circuit performs an adjoint operation of the data loader of a uniform norm-1 vector $\left(\frac{1}{\sqrt{r}},\frac{1}{\sqrt{r}},\ldots,\frac{1}{\sqrt{r}}\right)$, where $r=\max(m,n)$ represents the number of qubits excluding the ancillary qubit. This is followed by an X gate and a CNOT gate.
@@
Finally, we load $\left(\frac{1}{\sqrt{r}},\frac{1}{\sqrt{r}},\ldots,\frac{1}{\sqrt{r}}\right)$ and a Hadamard gate. In this way, the final quantum state in the circuit is
$\frac{1}{2}\sum_{j}\left(\sum_{i}W_{ji}x_{i}+\frac{1}{\sqrt{r}}\right)\ket{0,\mathbf{e}_{j}}+\frac{1}{2}\sum_{j}\left(\sum_{i}W_{ji}x_{i}-\frac{1}{\sqrt{r}}\right)\ket{1,\mathbf{e}_{j}},$ (1)
where |$\ket{\xi,\mathbf{e}_{j}}$ indicates the ancillary qubit is in state |$\ket{\xi}$ for $\xi\in\{0,1\}$, and the rest qubits are in the state |$\ket{\mathbf{e}_{j}}$. Thus, for each $y_{j}=\sum_{i}W_{ji}x_{i}$, the tomography includes two steps:
  1. 1.
    Sign recovery: noticing that
    ${\rm Pr}[0,\mathbf{e}_{j}]-{\rm Pr}[1,\mathbf{e}_{j}]=1/\sqrt{r}\sum_{i}W_{ji}x_{i}=y_{j}/\sqrt{r},$
    we have
    ${\rm sign}(y_{j})=\begin{cases}+1,\text{ if }{\rm Pr}[0,\mathbf{e}_{j}]\geq{\rm Pr}[1,\mathbf{e}_{j}]\\ -1,\text{ otherwise}\end{cases}.$
  2. 2.
    Value recovery:
    $y_{j}=\begin{cases}{\rm sign}(y_{j})\cdot\left(2\sqrt{{\rm Pr}[0,\mathbf{e}_{j}]}-\frac{1}{\sqrt{r}}\right),\text{ if }{\rm sign}(y_{j})>0\\ {\rm sign}(y_{j})\cdot\left(2\sqrt{{\rm Pr}[1,\mathbf{e}_{j}]}+\frac{1}{\sqrt{r}}\right),\text{ otherwise}\end{cases}.$ (2)
In this scheme, the error of the estimating $y_{j}$ is independent of the dimension of the problem $r$. As the tomography circuit has depth $\mathcal{O}(n)$, the tomography process maintains a complexity of $\mathcal{O}(n/\delta^{2})$ where $\delta$ is the tomography error and $1/\delta^{2}$ is the number of measurements. The scale of the tomography error is further explained in Section 4.1.
@@
When the output dimension is smaller than the input dimension ($m<n$), the tomography circuit is still the same, but only the information of bottom $m$ qubits are finally considered. In other words, the $\mathbf{e}_{j}$ in Eqs. (1) to (2) refers to $j$th unary state for the bottom $m$ qubits. Consequently, the output $\sum_{i}W_{ji}x_{i}$ is restricted to size $m$.
2.1.5 Error mitigation